Explain how you will teach the pythagoras theorem to grade 8 learners show all steps of your explanation
1
Expert's answer
2020-06-08T21:05:52-0400


We will get 4 right-angled triangle, hypotenuse of each of them is ‘a’: remaining sides of each of them are band c. Remaining part of the figure is the

square EFGH, each of whose side is a, so area of the square EFGH is a2.


Now, we are sure that square WXYZ = square EFGH + 4 ∆ GYF


or, (b + c)2 = a2 + 4 ∙ 1/2 b ∙ c


or, b2 + c2 + 2bc = a2 + 2bc


or, b2 + c2 = a2